OpenStudy (anonymous):

I am doing another integral.... I'll need a verification, if I am wrong, or when I am done.

OpenStudy (anonymous):

the problem is:\[\int\limits_{ }^{ }x \sin(x) \cos(x) dx\]

OpenStudy (anonymous):

I am thinking of differentiating the x, and integrating sin(x)cos(x).

OpenStudy (anonymous):

\[\int\limits_{ }^{ }\sin(x)\cos(x)~dx=-\sin^2(x)-\int\limits_{ }^{ }\cos^2(x)~dx\]

OpenStudy (anonymous):

just a tip. use the identity sin2x = 2sinxcosx =))

OpenStudy (anonymous):

oh, tnx, I'll definitely try that

OpenStudy (anonymous):

you still have question? feel free to ask :)

OpenStudy (anonymous):

\[\int\limits_{ }^{ }x \sin(x)\cos(x)~dx\]\[\frac{1}{2}\int\limits_{ }^{ }x \sin(2x)~dx\]

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

\[\frac{1}{2}\int\limits_{ }^{ }x \sin(2x)~dx=-\frac{1}{4}x \cos(2x)-\frac{1}{4}\int\limits_{ }^{ }\cos(2x)~dx\]

OpenStudy (anonymous):

\[\frac{1}{2}\int\limits_{ }^{ }x \sin(2x)~dx=-\frac{1}{4}x \cos(2x)+\frac{1}{8}\int\limits_{ }^{ }\sin(2x)+C\]

OpenStudy (watchmath):

the second term of your answer is incorrect.

OpenStudy (anonymous):

I have 1/2 from a chain of (2x), and it becomes positive, because inetgral of sin(x) is -cos(x). are you sure it's incorrect?

OpenStudy (anonymous):

I mena integral of cos(x) is sin(x). you are right it is -1/8

OpenStudy (watchmath):

the \(\frac18 \int \sin(2x)\,dx\) part

OpenStudy (anonymous):

\[\frac{1}{4}x \cos(2x)- \frac{1}{8} \sin(2x)+C\]

OpenStudy (watchmath):

great!

OpenStudy (anonymous):

so it's right?

OpenStudy (watchmath):

To make you confidence of your work, you can always take the derivative of your answer to verify that it is the same as the function that you integrate

OpenStudy (anonymous):

well, when I take the derivative I get. \[\frac{1}{4}\cos(2x)-\frac{1}{2}x \sin(2x)-\frac{1}{4}\cos(2x).\]\[-\frac{1}{2}x \sin(2x)\]\[-x \sin(x) \cos(x)\]

OpenStudy (anonymous):

I am getting a negative value....

OpenStudy (anonymous):

I went wrong on the derivative ...

OpenStudy (anonymous):

the answer should be:\[-\frac{1}{4}\cos(2x)-\frac{1}{8}\sin(2x)+C\]

OpenStudy (anonymous):

my bad, it's\[-\frac{1}{4}\cos(2x)-\frac{1}{8}x \sin(2x)+C\]

OpenStudy (anonymous):

now I will finally take the derivaitve.

OpenStudy (anonymous):

now, I am lost, I got the integral incorrectly, most likely.

OpenStudy (watchmath):

\(-\frac{1}{4}x\cos(2x)+\frac{1}{8}\sin(2x)+C\)

OpenStudy (anonymous):

wait it is positive 1/8 sin(2x) ?

OpenStudy (anonymous):

\[-\frac{1}{4}\cos(2x)+\frac{1}{8}\sin(2x)+C\]derivative.\[\frac{1}{2}x \sin(2x)-\frac{1}{4}\cos(2x)+\frac{1}{4}\cos(2x) \]

OpenStudy (anonymous):

which becomes,\[x \sin(x)\cos(x)\]